b) A news website reported (Feb. 17, 2020) that ridership of Karachi Metro Bus declined after a fare increase: “There were nearly four million fewer riders in December 2019, the first full month after the price of a token increased Rs. 25 to Rs.150, than in the previous December, a 4.3 percent decline.”
i. Use these data to estimate the price elasticity of demand for Karachi Metro Bus. (2 Marks)
ii. According to your estimate, what happens to the Karachi Metro Bus’s revenue when the fare rises? (0.5 Marks)
iii. Why might your estimate of the elasticity be unreliable?
1
Expert's answer
2020-11-30T16:42:46-0500

i) "\bold {Answer}"

"\eta_p =" "0.0086"


"\bold {Solution}"

We are given that quantity demanded declined by "4.3\%" in December 2019.

Price increased from Rs. 25 to Rs. 150


Percentage change in price

= "\dfrac {Rs. 150 - Rs. 25}{Rs. 25} × 100\%"


"= \dfrac {Rs. 125}{Rs. 25} × 100\%"


"= 500\%"


Now, "\eta_p = \dfrac {\% \space ∆Q}{\% \space ∆P}"


"= \dfrac {4.3\%}{500\%}"


"= 0.0086"


ii) "\bold {Answer}"

Revenue increases


"\bold {Explanation}"

The demand for Karachi Metro bus services is inelastic in demand because 0.0086 is less than 1. When demand is inelastic, an increase in price results in a less than proportionate decrease in quantity demanded. As a result, total revenue increases for every price increase.


iii) The estimate might be unreliable due to bias in estimation.

1. The decrease in demand as price increases might be seasonal. The estimation must be monitored throughout the year.


2. Other factors affecting demand, such as increase in competition and fall in income, must also be considered as they may cause a demand decrease.